OpenStudy (anonymous):

What is the percentage error if I use 3 1/7 for pi ?

OpenStudy (nikato):

http://www.algebra.com/algebra/homework/percentage.faq.question.209312.html U can check this out

OpenStudy (anonymous):

Thanks for the link,...but I honestly still don't get it~ Could you please explain it?

Parth (parthkohli):

Percentage error is in a way,\[\dfrac{|\text{new number}- \text{old number|}}{\text{old number}}\times 100\]

OpenStudy (anonymous):

The choice of numerator is sometimes unclear. Evaluate (100%)(3 + 1/7 - pi)/pi, where pi=3.14159..., available on your calculator.

OpenStudy (anonymous):

Okay, thanks! :)